Find f(ar) and f(z) for the following: cos(2rs) a) f (r) sin
Answer #1

a)   f(x) = exp(Co(2x3))

f ' (x) = exp(cos(2x3)) * (-sin(2x3)) * 6x = 6xSin(2x3) exp(Cos(2x3) )

f "(x) = 6 { xSin(2x3) * [6xSin(2x3) exp(Cos(2x3)] +

                        exp(Cos(2x3) [6Sin(2x3) + 6x* Cos(2x3) * 6x ] }

        = 6exp(Cos(2x3) { 6x2 Sin2(2x3) + 6 Sin(2x3) + 36x2 Cos(2x3) }

b) f(x) = 20x3 +12x2 -8xSin(7x)

f '(x) = 60 x2 +24x - 8 ( Sin(7x) +x * Cos(7x) *7)

          = 60 x2 +24x - 8sin(7x) -56xCos(7x)

f " (x) = 120 x + 24 - 8*7 *Cos(7x) - 56( Cos(7x) - x* Sin(7x) *7)

           = 120x +24 - 102 Cos(7x) +392 xSin(7x)